### v1

# Chinese Academy of Sciences, Beijing 100190, P. R. China

Abstract
Domain walls are topological defects produced by the spontaneous symmetry-breaking of discrete symmetry during cosmological phase transitions. The horizon-size domain wall can significantly contribute to the energy density in the late-evolution stage. We propose that the density perturbations from the fluctuations in the number density of the horizon-size domain wall could collapse to form primordial black holes. This mechanism becomes effective when the domain wall energy density ratio to that of the radiation reaches about 0.1 in the radiation-dominated universe. We find that models with $Z_{2}$ symmetry are excluded for interpreting pulsar timing array observations on the nano-Hz gravitational wave background since this model’s domain wall number density fluctuations could lead to an overabundance of the primordial black holes. Moreover, models with $N\sim 10$ domain walls also suffer strong constraints from the overabundance of primordial black holes.
1 Introduction
It is widely believed that new physics is required to explain various phenomena beyond the Standard Model (SM), including the neutrino mass and mixing King:2014nza , dark matter Bertone:2016nfn , strong CP problem Marsh:2015xka , and matter-antimatter asymmetry Dine:2003ax , etc. As the cosmic temperature drops with the expansion of the Universe, the spontaneous breakdown of discrete symmetries during a phase transition in new physics models can generate sheet-like topological defects, the domain walls Zeldovich:1974uw ; Kibble:1976sj . The Universe after the phase transition is then randomly divided by them into distinct domains, which represent the clumps of the degenerate vacua. The domain walls tend to stretch in size and the neighboring domains with the same vacua merge to stabilize the network. This evolution can be described by the percolation theory Stauffer:1978kr , whose fundamental results are based on the existence of a percolation threshold $p_{c}=0.311$ for a simple cubic lattice. The domain wall’s energy density scales as $\rho_{w}\propto t^{-1}$ in the scaling regime, which is much slower than the decay of radiation and dust, and eventually dominates the cosmological evolution. In the domain wall era, the repulsive force from domain walls would lead to an inflation phase of the Universe with the scale factor going as $a\propto t^{2}$. Such an expansion history is not compatible with the observations since it could strongly alter the Big Bang nucleosynthesis (BBN), reduce the time for the galaxy formation, and lead to unacceptably large anisotropies of the cosmic microwave background (CMB) Planck:2015fie . Nevertheless, the research on the nature of dark energy in the context of domain wall has been continuously conducted by recent works Friedland:2002qs ; Conversi:2004pi ; Mulki:2022tfw .
Discrete symmetries exist widely in new physics models. A $Z_{2}$ symmetry can emerge from the spontaneous breakdown of a gauge group in the grand unified theory (GUT) Preskill:1992ck ; Dunsky:2021tih . The $Z_{2}$ symmetry is also commonly imposed on the two-Higgs-doublet models for the spontaneous CP violation to suppress the Higgs-mediated flavor-changing neutral currents (FCNCs) at tree level Lee:1973iz ; Preskill:1991kd ; Gunion:2002zf . The global $U(1)$ symmetry in the axion models for solving the strong CP problem is explicitly broken to the $Z_{\mathcal{N}}$ symmetry by the QCD instanton effect Sikivie:1982qv ; Preskill:1991kd . Furthermore, to provide a dark matter candidate, a $Z_{2}$ symmetry is often imposed to make the dark matter particle stable Chao:2017vrq ; Huang:2017kzu ; Grzadkowski:2018nbc ; Chiang:2019oms ; Chiang:2020yym ; King:2023ayw ; Pham:2024vso .
A common solution to the domain wall problem is to introduce a bias potential $V_{\rm bias}$ to break the degenerate vacua Dine:1993yw ; Dvali:1994wv . The domain wall network collapses as its surface energy is comparable with the bias potential, $\sigma H\simeq V_{\rm bias}$. The bias potential can be generated by the instanton effects if the global symmetry is anomalous under the framework of the theory, with such examples as the QCD and quantum gravity Preskill:1991kd ; Dine:1993yw ; Dvali:1994wv ; Abel:1995wk ; Larsson:1996sp ; Lu:2023mcz . It is notable that based on the successful and precise prediction in Ref. Chiang:2020aui , Ref. Lu:2023mcz suggest that the recent pulsar timing array (PTA) observations, including the NANOGrav 15-year dataset (NG15) and IPTA-DR2 dataset, strongly support the mechanism of the QCD instanton-induced domain wall annihilations.
Primordial black holes (PBHs) have attracted a lot of attention since it was first suggested by Zel’dovich more than 50 years ago Zeldovich:1967lct . Contrary to the black holes resulting from the death and collapse of stars in their final evolution stage, PBHs are formed by cosmological processes in the early Universe. In 1971, Hawking proposed the first mechanism of PBH production Hawking:1971ei ; Carr:1974nx : the fluctuations of a radiation-dominated Universe can produce some overdense regions. Once the density of the region $\delta=\Delta\rho/\bar{\rho}$ exceeds a critical value $\delta_{c}\simeq w$, the overdense region can stop expanding and re-collapse to form PBHs when the region enters the horizon. Cosmic inflation Ivanov:1994pa ; Randall:1995dj ; Garcia-Bellido:1996mdl is commonly considered a primary source of density perturbation. The large curvature perturbations generated by the quantum fluctuations of the scalar fields are dominated by the density perturbations once they exit the horizon Garcia-Bellido:1996mdl . Other sources for PBH production include the collapse of topological defects such as cosmic strings Hawking:1987bn ; Polnarev:1988dh ; Brandenberger:2021zvn and closed domain walls Ipser:1983db ; Ferrer:2018uiu ; Liu:2019lul ; Gouttenoire:2023gbn ; Ge:2023rrq , the density perturbation from Q-balls during a matter-dominated era Cotner:2016cvr ; Cotner:2019ykd , the collision of bubble walls Crawford:1982yz ; Kodama:1982sf ; Hawking:1982ga ; Johnson:2011wt ; Kusenko:2020pcg and the collapse of false-vacuum patches during a first-order phase transition Baker:2021sno ; Gross:2021qgx ; Kawana:2021tde ; Liu:2021svg ; Baldes:2023rqv ; Flores:2024lng ; Gouttenoire:2023naa ; Lewicki:2023ioy .
In contrast to PBH formation from the collapse of closed domain wall Ipser:1983db ; Ferrer:2018uiu ; Liu:2019lul ; Gouttenoire:2023gbn ; Ge:2023rrq , we show in this work that the density perturbation due to the number density fluctuations of horizon-size domain walls in the late stage could lead to the formation of PBHs. The collapse of a closed domain wall needs to satisfy the Schwarzschild criterion, i.e., the closed wall has a Schwarzschild radius near collapse Ferrer:2018uiu . Our scenario is based on the critical collapse of horizon-size perturbations, as outlined above. We find that if the annihilations of the domain wall are the source of the nano-Hz stochastic gravitational wave background (SGWB) observed recently by PTAs, a significant PBH relic abundance could be produced via the number density fluctuations of the walls before their annihilations.
This work is arranged as follows. In section 2, we briefly review the equation of state and evolution of the domain wall. In section 3, we calculate the variance of density contrast from domain wall number density fluctuations. In section 4, we review the critical collapse mechanism. In section 5, we calculate the PBH relic abundance from domain wall number density fluctuations. In section 6, we define an effective equation of state for the mixture gas containing domain walls and radiation. A fit formula for the $\delta_{c}-\omega_{\rm eff}$ relation from the simulation is also provided. We derive the gravitational wave background from the domain wall annihilations in section 7, and perform the Bayesian statistics analysis of the IPTA-DR2 and NG15 datasets in section 8. In section 9, We show the constraints on the domain wall interpretations for the PTAs using the PBH abundance from the domain wall number density fluctuations. Finally, in section 10, we summarize our conclusions and discuss the uncertainty in the PBH formations. Various particle physics models for the realization of $N$ domain walls for PTAs are also discussed in this section. In appendix A, appendix B, and appendix C, we provide respectively the details of calculations for the energy-momentum tensor of the domain wall in a $Z_{2}$ symmetry model, the surface energy tensor in the observer frame, and the velocity-dependent one-scale (VOS) model. In appendix D, we provide details for the complementary error function. In appendix E, we present further details of our numerical strategy for the PTA dataset.
2 Domain wall
In this section, we first calculate the statistical state of the domain walls. To make sense of the domain wall evolution, we review the VOS model with the consideration of particle friction. The conclusions from this model are generally compatible with numerical simulations.
2.1 Domain wall equation of state
The energy-momentum tensor of a plane domain wall lying in the $yz$ plane at $x=0$ (neglecting the thickness of the wall) can be estimated as Vilenkin:1984ib ; Gelmini:1988sf
$T_{\mu\nu}(x)=\sigma_{w}\delta(x){\rm diag}(1,0,-1,-1),$ (1)
where $\sigma_{w}$ is the surface tension of the wall. In appendix A, we provide a concrete example for the calculation of the energy-momentum tensor in a toy model with a $Z_{2}$ symmetry. Consider that within a volume $V\equiv L^{3}$, there are $N$ $yz$-plane walls moving along the $x$ direction with an average velocity $\overline{v}_{w}$. The average energy-momentum tensor within a length $L$ can be estimated as
[…]
where $\overline{L}=L/N$ is the average wall separation, in which there exists on the average one domain wall with the surface energy-momentum tensor
[…]
with the superscript $x$ denoting the axis of wall motion. The energy-momentum tensor (1) is in the rest frame of the wall. We can obtain the energy-momentum tensor in the observer frame by performing a Lorentz transformation. In general, the domain wall can move along an arbitrary direction in the three-dimensional space and, with some algebra, we obtain the surface energy-momentum tensor in the observer frame as
$S_{\mu\nu}(\overline{v}_{w})=\frac{\sigma_{w}}{3}{\rm diag}(\gamma^{2},\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2),$ (4)
where $\gamma=1/\sqrt{1-\overline{v}_{w}^{2}}$ is the Lorentz factor. See appendix B for the details of calculations.
The average energy-momentum tensor is determined by $\overline{T}_{\mu\nu}=S_{\mu\nu}/\overline{L}$. Note that since the wall can extend along any direction, we interpret $\overline{L}$ as the correlation length of the wall. The energy density and pressure of the wall are given by temporal and spatial components of the energy-momentum tensor
$\rho_{w}=\overline{T}_{00}=\frac{\gamma^{2}\sigma_{w}}{\overline{L}},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ p_{w}=\overline{T}_{ii}=\frac{\sigma_{w}}{3\overline{L}}(\overline{v}_{w}^{2}\gamma^{2}-2).$ (5)
We then obtain the equation of state of the wall
$\omega=\frac{p_{w}}{\rho_{w}}=\frac{-2+\overline{v}_{w}^{2}\gamma^{2}}{3\gamma^{2}}.$ (6)
For a relativistic wall with velocity $\overline{v}_{w}\sim 1$ and $\gamma\gg 1$, we have $\omega\simeq 1/3$, and the wall behaves like radiation. For a static wall, $\overline{v}_{w}\simeq 0$ and $\omega\simeq-2/3$. The equation of state is negative, $\omega<0$, for $\overline{v}_{w}\lesssim 0.8$. In this case, the wall contributes negative pressure to the energy-momentum tensor and can lead to the inflation of the universe if it dominates the energy density of the universe.
2.2 The evolution of a domain wall
The correlation length $\overline{L}$ and velocity $\overline{v}_{w}$ are two main physics quantities for describing the domain walls. The evolution of $\overline{L}$ and $\overline{v}_{w}$ with the cosmic expansion can be described by the VOS model Martins:2016book (see appendix C for more details)
$\frac{d\overline{L}}{dt}$ $=H\overline{L}+\frac{\overline{L}}{\ell_{d}}\overline{v}_{w}^{2}+c_{w}\overline{v}_{w},$ (7)
$\frac{d\overline{v}_{w}}{dt}$ $=\left(1-\overline{v}_{w}^{2}\right)\left(\frac{k_{w}}{\overline{L}}-\frac{\overline{v}_{w}}{\ell_{d}}\right),$
where the momentum parameter $k_{w}=0.66\pm 0.04$ and the chopping parameter $c_{w}=0.81\pm 0.04$ from the simulations of domain wall evolution in a radiation-dominated universe Martins:2016ois . The damping length is defined as
$\frac{1}{\ell_{d}}=3H+\frac{1}{\ell_{f}}.$ (8)
The first term on the right-hand side of Eq. (8) describes the Hubble damping effect due to the cosmic expansion and the second term denotes the particle friction, which is related to the particle pressure by $\ell_{f}=\overline{v}_{w}\sigma_{w}/\Delta P$, where the pressure between the wall and particles $\Delta P$ is model-dependent.
Figure 1: The evolution of $\overline{L}$ (red curves) and $\overline{v}_{w}$ (blue curves) as a function of cosmic time $t/t_{s}$. The results without any particle friction are denoted by the solid lines. From the upper left to the lower right, the dot-dashed lines represent the results with $N_{c}=1$, 5, 15, and 30, respectively.
In Fig. 1, we depict the evolution of the domain wall formed at time $t_{s}$ with a symmetry-breaking scale $v=\sigma_{w}^{1/3}=10^{5}$ GeV. We also assume a pressure induced by friction from the interaction between wall and QCD gluons given by
$\Delta P=\frac{24N_{c}^{2}\alpha_{s}^{2}}{\pi^{4}}\overline{v}_{w}T^{4},$ (9)
where we will adopt the strong coupling $\alpha_{s}=1$ for illustration and note that $\alpha_{s}\lesssim 0.1$ for the energy scale at $\sim 10^{5}$ GeV, and the strength of the pressure is controlled by the parameter $N_{c}$. For the solid curves, the friction is removed by setting $N_{c}=0$. From the upper left panel to the lower right panel of Fig. 1, the dot-dashed curves represent the results with $N_{c}=1$, 5, 15, and 30, respectively. The results without particle friction indicate that after its formation, the domain wall rapidly expands to a superhorizon size, $\overline{L}\sim 6t$, due to the tension of the wall. At time $t\sim 20t_{s}$, the Hubble friction from the cosmic expansion becomes significant and damps the further expansion of the wall. The evolution of the domain wall finally enters into the scaling regime, in which the domain wall is of the horizon size, $\overline{L}\simeq t$, and the velocity of the wall $\overline{v}_{w}\simeq 0.4$. For common models with $N_{c}=1$, the friction effect from interaction with gluons is negligible. When $N_{c}\gtrsim 5$ the suppression from particle friction on the domain wall growing at $t\lesssim 20t_{s}$ becomes significant. The suppression effect stops improving much for $N_{c}\gtrsim 15$. A nontrivial conclusion from Fig. 1 is that the sizable friction (if existing in the model) from the surrounding plasma can not stop the walls from eventually entering into the scaling regime (for $t\gtrsim 10^{4}t_{s}$). Therefore, the mechanism of the PBH production from domain wall number density fluctuations is independent of the particle physics models.
3 Density fluctuations from domain walls
After the production of domain walls via the Kibble mechanism Kibble:1976sj , the size of the wall quickly grows to the scale of Hubble horizon $r_{H}=1/H$. Therefore, the domain wall network is a few-body system, which can lead to large density fluctuations. Consider a network consisting of $N$ identical walls with mass $M_{w}$ in a Hubble volume $V_{H}\equiv 4\pi r_{H}^{3}/3$ (this assumption is reasonable at least for domain walls in the scaling regime, where domain walls have a similar dimension $\overline{L}\simeq t$). We employ the Poisson distribution to describe the probability of finding $k$ walls in a volume $V$
$\mathbb{P}(k,V)=\frac{(\bar{n}V)^{k}}{k!}e^{-(\bar{n}V)},$ (10)
where $\bar{n}=N/V_{H}$ is the average number density over the Hubble horizon volume $V_{H}$ (or the background number density). As expected, the ensemble average of the number density $n=k/V$ in a volume $V$ is equal to the background number density,
$\langle n\rangle=\sum_{k=0}^{\infty}n\mathbb{P}(k,V)=\sum_{k=0}^{\infty}\left(\frac{k}{V}\right)\mathbb{P}(k,V)=\bar{n}.$ (11)
The mean square number density fluctuations or the variance from the background is given by
$\langle(\Delta n)^{2}\rangle=\langle n^{2}\rangle-\bar{n}^{2}=\sum_{k=0}^{\infty}\left(\frac{k}{V}\right)^{2}\mathbb{P}(k,V)-\bar{n}^{2}=\frac{\bar{n}}{V}.$ (12)
Let $\delta=\Delta\rho_{w}/\bar{\rho}_{\rm tot}$ be the relative energy density perturbation (density contrast), where the total background energy density $\bar{\rho}_{\rm tot}=\bar{\rho}_{w}+\bar{\rho}_{\rm rad}$, with $\bar{\rho}_{\rm rad}$ being the radiation background energy density. We then define the relative mean square energy density fluctuations (the variance of density perturbation) due to the Poisson fluctuations in the volume $V$ as
$\sqrt{\langle\delta^{2}\rangle}=\frac{\sqrt{\langle(\Delta\rho_{w})^{2}\rangle}}{\bar{\rho}_{\rm tot}}=\frac{\sqrt{\langle(\Delta\rho_{w})^{2}\rangle}}{\bar{\rho}_{w}}\frac{\bar{\rho}_{w}}{\bar{\rho}_{\rm tot}}.$ (13)
Assuming that the domain wall is the only source for the density fluctuation, i.e., $\rho_{\rm rad}=\bar{\rho}_{\rm rad}$, we have $\Delta\rho_{w}=\rho_{w}+\rho_{\rm rad}-\bar{\rho}_{\rm tot}=\rho_{w}-\bar{\rho}_{w}$. The density fluctuations can be written as
$\langle(\Delta\rho_{w})^{2}\rangle=\langle\rho_{w}^{2}\rangle-\bar{\rho}_{w}^{2}=M_{w}^{2}\langle(\Delta n)^{2}\rangle$ (14)
where $\rho_{w}=nM_{w}$. Then the relative density fluctuations of the total density in a volume $V$ are given by
$\sqrt{\langle\delta^{2}\rangle}=\frac{1}{\sqrt{\bar{n}V}}\frac{f_{w}}{f_{w}+1}\simeq\frac{f_{w}}{\sqrt{\bar{n}V}},$ (15)
where the energy density fraction of the domain wall $f_{w}=\bar{\rho}_{w}/\bar{\rho}_{\rm rad}\ll 1$ in a radiation-dominated universe.
Then the probability density distribution function of the density contrast, $P(\delta)$, is featured by a zero ensemble averaged value, $\langle\delta\rangle=0$, and a variance $\sigma$ given by Eq. (15). In this work we consider a Gaussian density distribution function
$P(\delta)=\frac{1}{\sqrt{2\pi}\sigma}\exp\left(-\frac{\delta^{2}}{2\sigma^{2}}\right).$ (16)
We consider the horizon-size density perturbation induced by the domain walls in the scaling regime, the variance of the probability distribution is given by
$\sigma=\left.\sqrt{\langle\delta_{w}^{2}\rangle}\right|_{V\to V_{H}}=\frac{f_{w}}{\sqrt{\bar{n}V_{H}}}=\frac{f_{w}}{\sqrt{N}}.$ (17)
Simulations Press:1989yh ; Hiramatsu:2012sc ; Lazanu:2015fua indicate that the domain walls tend to merge with the neighboring ones in the evolution, resulting in a wall network with a small number of horizon-size walls. For a model with a $Z_{2}$ symmetry, there eventually exists $N=1$ domain wall that separates two domains Press:1989yh . For models with $\mathcal{N}\sim 10$ domains, one expects that $N\sim\mathcal{N}$ domain walls exist in the late stage of the network evolution. Therefore, once the domain walls contribute significantly to the total energy density, $f_{w}\sim 0.1$, the energy density perturbation due to the number density fluctuations of the domain wall can be remarkable.
4 Critical collapse
Let us assume that the perturbed region is spherically symmetric and has an initial radius $R(t)$ larger than the Hubble scale. The locally overdense region can then constitute a spatially closed universe with a metric given by Carr:1975qj ; Yokoyama:1995ex
$ds^{2}=-dt^{\prime 2}+R^{2}\left(t^{\prime}\right)\left[\frac{dr^{2}}{1-\kappa r^{2}}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right)\right],$ (18)
where $\kappa>0$ denotes the perturbed total energy per unit mass. Therefore, the density of the perturbed region is
$\rho=\frac{3}{8\pi G}\left(H^{2}+\frac{\kappa}{R^{2}}\right).$ (19)
The initial expansion rates of the background and the perturbed region can be set to be the same by choosing the coordinates Harrison:1969fb ; Yokoyama:1995ex so that the initial density contrast is given by
$\delta_{i}=\frac{\rho_{i}-\bar{\rho}_{i}}{\bar{\rho}_{i}}=\frac{3\kappa}{8\pi G\bar{\rho}_{i}R_{i}^{2}}=\frac{\kappa}{H_{i}^{2}R_{i}^{2}},$ (20)
where the subscript $i$ is used to denote the initial time $t_{i}$ and the initial background density $\bar{\rho}_{i}=3H_{i}^{2}/(8\pi G)$. Integrating the continuity equation
$\frac{d\rho}{dt}+3H(\rho+p)=0,$ (21)
we obtain the evolution of the energy density
$\rho=\rho_{i}\left(\frac{R}{R_{i}}\right)^{-3(1+\omega)},$ (22)
with a constant equation of state $\omega\equiv p/\rho$. At a critical time $t_{c}$, the overdense regions stop expanding, i.e. $H(t_{c})=0$, with a critical density $\rho_{c}=\kappa/R_{c}^{2}$. The ratio between the energy density at the initial time and the critical time is then given by
$\frac{\rho_{c}}{\rho_{i}}=\frac{\kappa}{R_{i}^{2}H_{i}^{2}+\kappa}\frac{R_{i}^{2}}{R_{c}^{2}}=\frac{\delta_{i}}{1+\delta_{i}}\frac{R_{i}^{2}}{R_{c}^{2}}.$ (23)
Together with Eqs. (22) and (23), we obtain the critical scale factor that the overdense region stops expanding
$\frac{R_{c}}{R_{i}}=\left(\frac{1+\delta_{i}}{\delta_{i}}\right)^{\frac{1}{1+3\omega}}.$ (24)
The scale factor evolves as $R(t)\propto t^{2/3(1+\omega)}$, therefore the critical time is given by
$\frac{t_{c}}{t_{i}}=\left(\frac{1+\delta_{i}}{\delta_{i}}\right)^{\frac{3(1+\omega)}{2(1+3\omega)}}.$ (25)
The overdense region could contrast further against the pressure gradient if the self-gravitation $\Phi$ of the mass within the critical radius $R_{c}$ is larger than the internal energy $U$ of the overdense region, which gives Carr:1974nx
$\Phi+U\sim-G\rho^{2}R_{c}^{5}+\omega\rho R_{c}^{3}\gtrsim 0.$ (26)
For a radiation-dominated universe, $\rho\sim t^{2}/G$, then one obtains the lower limit on the scale of the overdense region
$R_{c}\gtrsim R_{J}\simeq c_{s}t_{c},$ (27)
where $R_{J}$ is the Jeans scale and $c_{s}\simeq\sqrt{\omega}$. With Eqs. (24), (25), and (27), one finds the critical density criterion for PBH formation in the radiation-dominated universe:
$\delta_{i}\gtrsim\delta_{c}\simeq c_{s}^{2}\simeq\frac{1}{3}.$ (28)
A more precise value is $\delta_{c}=0.453$ for $\omega=1/3$ from the numerical calculation (see section 6). If the overdense region stops expanding at a radius larger than the Hubble horizon, i.e., $R_{c}\gtrsim t_{c}$, the PBHs formed in this region would be separated from our universe and, thus, make no contributions to the observations. We therefore have the limit on the size of the overdense region, $c_{s}t_{c}\lesssim R_{c}\lesssim t_{c}$, which corresponds to the condition
$\frac{1}{3}\lesssim\delta_{i}\lesssim 1.$ (29)
Since the density perturbations are induced by the number density fluctuations of the domain walls, the scale of the density perturbation should have a characteristic size as that of the walls and the evolution of the perturbation should also follow the evolution of the walls. As shown in section 2.2, the domain wall in the scaling regime with density $\rho_{w}\simeq\sigma_{w}/t$ has a horizon-size characteristic length $\overline{L}\simeq t$. The number density fluctuations of the wall during radiation domination would generate horizon-size density fluctuations, whose average amplitude scales with time as
[…]
A rough estimate using $\delta\sim\rho_{w}/\rho_{\rm rad}\sim G\sigma_{w}t$ would lead to the same conclusion. Domain walls are continuous sources for inducing density perturbations. At a given time, the density fluctuations are randomly generated by the number density fluctuation. Equation (30) indicates that the probability for the production of a large-density contrast grows with the increase of cosmic time. Those perturbations with amplitudes that exceed the critical value $\delta_{c}$ might undergo critical collapse and lead to PBH formation.
We note in passing that a recent study Flores:2024lng argued that the critical density criterion could not be applied to horizon-size density perturbations in the inflationary scenarios since the over-density would be generated by the curvature perturbation, without which the horizon-size patch had only expansionary solutions. We note that the condition of having a superhorizon-size perturbation is the result of inflation, rather than the requirement for the critical collapse. Obviously, the domain wall scenario we consider in this work is not inflation-based and does not suffer from such a problem. In fact, the initial estimate by Hawking and Carr Carr:1974nx required merely a horizon-size perturbation for the critical collapse.
5 PBH formation
We first calculate the PBH relic abundance from the critical collapse. We then review the power spectrum of the density perturbation from the inflation and summarize the constraints of current CMB observations on the inflationary scenarios. Finally, we calculate the variance of the density contrast from the number density fluctuations of the domain walls.
5.1 PBH relic abundance
Let us define a parameter $\beta$ to represent the fraction of total energy consisting of PBHs at their formation time $t_{f}$:
$\beta(t_{f})=\frac{\rho_{\mathrm{PBH}}(t_{f})}{\rho_{\rm tot}(t_{f})}=\Omega_{\mathrm{PBH}}\left(\frac{H_{0}}{H_{f}}\right)^{2}\left(\frac{a_{0}}{a_{f}}\right)^{3},$ (31)
where $\Omega_{\mathrm{PBH}}\equiv\rho_{\mathrm{PBH}}\left(t_{0}\right)/\rho_{\text{crit}}$ (with $\rho_{\rm crit}=3H_{0}^{2}/(8\pi G)$) is the density of the PBH relic abundance at present and 00 denotes the present time. In the second equality of Eq. (31) we have used $\rho_{\rm PBH}(t_{f})=\rho_{\rm PBH}(t_{0})(a_{0}/a_{f})^{3}$. It is commonly related to the dark matter relic abundance by $\Omega_{\mathrm{PBH}}=f_{\mathrm{PBH}}\Omega_{\mathrm{CDM}}$, with the fraction of dark matter $f_{\rm PBH}$ in the form of PBHs.
It is expected that the PBHs formed from the Hubble horizon (re-)entry of the large density perturbations could have masses at the same order as the associated horizon mass $M_{H}\simeq 2\times 10^{15}(t/1s)M_{\odot}$. To further estimate the PBH production, let’s take the naive relation between the PBH mass $M$ and the horizon mass from critical phenomena Sasaki:2018dmp ; Carr:2020gox
$M=\gamma_{p}M_{H},$ (32)
where the numerical factor $\gamma_{p}\simeq 0.2$. During the radiation domination, the expansion rate of the universe is $H=1/(2t)$, and the energy density of a flat universe is
$\rho_{\rm rad}=\frac{3H^{2}}{8\pi G}=\frac{\pi^{2}}{30}g_{*}(T)T^{4},$ (33)
where $g_{*}(T)$ is the total effective degrees of freedom, and the Newtonian gravitational constant $G$ is related to the Planck mass $M_{\rm pl}$ by $G=1/M_{\rm pl}^{2}$ with $M_{\rm pl}=1.22\times 10^{19}$ GeV. The horizon mass is then given by
$M_{H}=\frac{4\pi}{3}\rho_{\rm rad}r_{H}^{3}=\frac{1}{2GH}=\frac{3\sqrt{5}M_{\rm pl}^{3}}{4\pi^{3/2}g_{*}^{1/2}(T)T^{2}}.$ (34)
Taking advantage of the Press-Schechter formalism, the PBH energy fraction can be expressed as Niemeyer:1997mt
[…]
where we consider a Gaussian probability distribution (16) for the density contrast. Note that we use an upper limit of infinity for the integration and confirm that it is not different from the integration result with an upper limit of $\delta=1$ from the requirement of avoiding separate universes Carr:1974nx ; Yokoyama:1995ex . A recent work Kopp:2010sh argued that this upper limit was a result of the geometry of the selected slicing rather than as a consequence of avoiding separate universes, which instead led to a divergence of density fluctuations on the superhorizon scale. With the PBH mass given by Eq. (32), the fraction of PBHs at the formation time is
$\beta(M)=\gamma_{p}{\rm erfc}\left(\frac{\delta_{c}}{\sqrt{2}\sigma}\right),$ (36)
where ${\rm erfc}(x)$ is the complementary error function (see appendix D for more details). The critical density contrast $\delta_{c}$ varies with the equation of state $\omega$. For radiation phase $\delta_{c}=0.453$. We determine $\delta_{c}$ with the $\delta_{c}-\omega$ relation revealed by the simulations in Ref. Musco:2012au (see section 6).
A more precise PBH mass formation was obtained from the numerical computations on the collapse of a relativistic fluid, which indicated a simple scaling law for the PBH mass distribution function near the critical threshold Choptuik:1992jv ; Evans:1994pj ; Koike:1995jm
$M=CM_{H}(\delta-\delta_{c})^{b},$ (37)
where the constant parameter $C=3.3$ and $b=0.36$ for the formation of PBH during radiation domination Byrnes:2018clq . Note that these parameters may be affected by the phase of cosmic matter. Then the PBH energy fraction is given by
[…]
In this work, we use the PBH mass (32) by default, unless otherwise stated.
Once the energy fraction in PBHs at their formation is obtained, the relic abundance of PBHs today can be determined by taking into account the cosmic expansion (and neglecting the evaporation and accretion of PBHs). The radiation density red-shifts as $a^{-4}$ while the PBH density scales as $a^{-3}$. Using the fact of entropy conservation in the comoving volume, i.e.,
$s_{f}a_{f}^{3}=s_{0}a_{0}^{3},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ \leavevmode\nobreak\ s(T)=\frac{2\pi^{2}}{45}g_{*,s}T^{3},$ (39)
where the current universe entropy density $s_{0}=2891.2\leavevmode\nobreak\ \rm cm^{-3}$. Together with Eqs. (31)-(33) we obtain the the current PBH relic density abundance
$\Omega_{\mathrm{PBH}}$ $=\frac{2\pi G\beta(M)s_{0}g_{*}(T)T}{H_{0}^{2}g_{*,s}(T)}=(3\sqrt{5\pi})^{1/2}\frac{\beta(M)s_{0}}{H_{0}^{2}(M_{\rm pl}M_{H})^{1/2}}\frac{g_{*}^{3/4}(T)}{g_{*,s}(T)}$ (40)
$\simeq\left(\frac{\beta(M)}{5.74\times 10^{-9}}\right)\left(\frac{0.68}{h}\right)^{2}\left(\frac{g_{*}(T)}{10.75}\right)^{3/4}\left(\frac{10.75}{g_{*,s}(T)}\right)\left(\frac{M_{\odot}}{M_{H}}\right)^{1/2},$
where the Hubble expansion rate is $H_{0}=100h\leavevmode\nobreak\ \mathrm{km\leavevmode\nobreak\ s}^{-1}\leavevmode\nobreak\ \mathrm{Mpc}^{-1}$ with $h=0.68$ from the Planck observation Planck:2018vyg . With the relation $M=\gamma_{p}M_{H}$ and the approximation $g_{*}(T)\simeq g_{*,s}(T)$, we reproduce $\Omega_{\rm PBH}$ given in Ref. Carr:2009jm . We observe that the PBH energy fraction at the formation should be of order $\beta(M)\sim 10^{-9}$ so that the PBH relic abundance today does not exceed the required dark matter relic abundance. Thus the variance of the density contrast should be small and the probability density distribution should be concentrated so that the probability for the production of large density contrasts is exponentially suppressed in the tail of the probability density distribution.
5.2 Inflation-induced density perturbation
From the above, we know that one of the most important parameters to determine the PBH abundance is the variance $\sigma$, which is directly related to the physics of the production of density perturbations.
Let’s first briefly review the PBH production in the inflation scenario. For the density fluctuations from the curvature perturbations during inflation, the variance can be estimated as
$\sigma^{2}=\int_{0}^{\infty}W^{2}(kR)\mathcal{P}_{\delta}(k)\frac{dk}{k},$ (41)
where $W(kR)$ is a volume-normalized Fourier transform of the window function smoothed over a comoving scale $R$ and can be assumed as Gaussian. On comoving hypersurfaces, the density perturbation can be related to the curvature perturbation by Green:2004wb
$\mathcal{P}_{\delta}(k)=\frac{4(1+\omega)^{2}}{(5+3\omega)^{2}}\left(\frac{k}{aH}\right)^{2}\mathcal{P}_{\mathcal{R}_{c}}(k).$ (42)
The PBHs are formed at the horizon crossing $k=aH$. The power-law ansatz is commonly adopted for the primordial curvature power spectrum:
$\mathcal{P}_{\mathcal{R}_{c}}(k)=A_{\mathcal{R}_{c}}\left(\frac{k}{k_{*}}\right)^{n_{s}-1},$ (43)
where the amplitude $A_{\mathcal{R}_{c}}$ is a constant that depends on the inflation physics, and $k_{*}$ is the reference scale. Current CMB and LSS observations on large scales ($k\lesssim{\rm Mpc^{-1}}$) have restricted the primordial curvature power spectrum at the order around $10^{-9}$. To have a sizable PBH formation for LIGO observation, the primordial curvature power spectrum should be of order $\mathcal{P}_{\mathcal{R}_{c}}(k)\sim 10^{-2}-10^{-1}$ (corresponding to a variance in the range $10^{-4}\lesssim\sigma^{2}\lesssim 10^{-2}$ which gives $\beta\sim 10^{-9}-10^{-6}$, see Fig. 10 in appendix D), which requires a blue-tilted spectral index $n_{s}\simeq 1.85$ at scale $k\sim 10^{6}\leavevmode\nobreak\ \rm Mpc^{-1}$ (corresponding to a PBH mass $M\sim 30M_{\odot}$asaki:2018dmp ; Byrnes:2018clq . On the other hand, the CMB observations have constrained the spectral index to be red-tilted, $n_{s}=0.968\pm 0.006$ at the scale $k_{*}=0.05\leavevmode\nobreak\ \rm Mpc^{-1}$. The relatively large blue-tilted spectral index at a small scale cannot be realized in the single-field slow-roll inflation but may be achieved in the running mass inflation or multi-scalar inflation models, in which the scale-dependent spectral index can be realized (see Refs. Stewart:1997wg ; Drees:2011hb ; Alabidi:2012ex ; Silk:1986vc ; Randall:1995dj ; GarciaBellido:1996mdl ; Clesse:2015wea for example).
5.3 Domain wall induced density perturbation
Compared with the inflation during which the density fluctuations are only induced at the surface of the last scattering, the metric perturbations are continuously generated by the domain wall network in the scaling regime. The domain wall tends to stretch in size and concatenate the nearby walls to form larger domains. This results in a domain wall network with large wall sizes and a small wall number at the late evolution stage. Therefore, one expects that the density perturbation and, thus, the power spectrum induced by the domain walls are most significant at the later stage. Such a picture is confirmed by recent simulations of the domain wall evolution to calculate the power spectrum Lazanu:2015fua . A simple estimate indicates that the power spectrum of the domain walls should have $\mathcal{P}_{w}\sim(G\sigma_{w}t)^{2}$, which is consistent with the variance given by Eq. (17). Thus, the density perturbations induced by the few-body system consist of domain walls that grow with time even in the radiation-dominated universe.
As has been shown in section 2.2 as well as the simulations, the energy density of the domain wall in the scaling regime is given by
$\rho_{w}=\frac{\mathcal{A}\sigma_{w}}{t}=2\mathcal{A}H\sigma_{w},$ (44)
where $\mathcal{A}$ depends on the number of walls. For $N\leq 6$ we take the values of $\mathcal{A}$ from the simulations Kawasaki:2014sqa , and for $N>6$ we take $\mathcal{A}=0.4N$. Therefore, the energy fraction of the wall is
$f_{w}=\frac{\rho_{w}}{\rho_{\rm rad}}=\frac{16\pi G\mathcal{A}\sigma_{w}}{3H}=\frac{32\pi G\mathcal{A}\sigma_{w}t}{3}.$ (45)
The variance of density contrast induced by the domain wall is then given by
$\sigma$ $=$ $\frac{f_{w}}{\sqrt{N}}=\frac{32\pi G\mathcal{A}\sigma_{w}t}{3\sqrt{N}}=\frac{32\pi G^{2}M_{H}\mathcal{A}\sigma_{w}}{3\sqrt{N}}$ (46)
$\simeq$ $1.81\times 10^{-3}\frac{\mathcal{A}}{\sqrt{N}}\frac{M_{H}}{M_{\odot}}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ \rm GeV}\right)^{3}.$
To make sure that the energy density of the domain wall does not dominate the universe we require $f_{w}\lesssim 1$, which gives the following constraint on the PBH horizon mass
$M_{H}\lesssim 5.52\times 10^{2}M_{\odot}\left(\frac{10^{5}\leavevmode\nobreak\ \rm GeV}{\sigma_{w}^{1/3}}\right)^{3}.$ (47)
Figure 2: Left: The variance $\sigma$ as a function of the PBH mass $M$. Right: The PBH relic abundance $\Omega_{\rm PBH}$ as a function of the PBH mass $M$. We take the wall number $N=1$.
In Fig. 2, we plot the variance (46) and PBH relic abundance as a function of the PBH mass with the relation $M=\gamma_{p}M_{H}$ and assuming the domain wall number $N=1$. The blue and green solid curves represent the results with the wall surface tension $\sigma_{w}^{1/3}=10^{5}$ GeV and $10^{6}$ GeV, respectively. As shown in the figure, the PBH relic abundance from the domain wall number fluctuations becomes significant when the variance $\sigma(=f_{w}/\sqrt{N})\sim 10^{-1}$, which would correspond to a PBH mass of $M\sim 10^{1}\leavevmode\nobreak\ M_{\odot}$ and $10^{-2}\leavevmode\nobreak\ M_{\odot}$ for $\sigma_{w}^{1/3}=10^{5}$ GeV and $10^{6}$ GeV, respectively, as shown in the right panel of the figure. The PBH abundance sharply increases with the increasing of domain wall energy fraction $f_{w}$ and reaches its maximum at $f_{w}=1$. For $f_{w}\simeq 1$, the PBH energy fraction reaches its maximum $\beta(M)\simeq 2\times 10^{-1}$ (see the right plot of Fig. 10 in appendix D). Beyond the maximum point, $\Omega_{\rm PBH}$ scales as $M^{-1/2}$. However, we note that the results with $f_{w}\gtrsim 1$ are not reliable. This is because the domain wall energy density dominates the radiation and the universe enters into a solid-like phase with the equation of state $\omega\simeq-2/3$. In this phase, the universe would undergo an inflation stage before the annihilation of the domain walls. In this work, we assume a radiation-dominated universe with $f_{w}\lesssim 1$ and, therefore, do not give a further discussion on this scenario.
6 Effective equation of state
The critical collapse scenario states that the horizon-size perturbations could stop expanding and collapse to form PBHs when they enter into the horizon with a density contrast exceeding a critical threshold $\delta_{c}$. Therefore, the exact value of the critical density contrast $\delta_{c}$, which depends on the equation of state $\omega$, is nontrivial for the PBH formation. Effects of $\omega$ on PBH formation have been investigated in the context of first-order phase transition Jedamzik:1999am ; Jedamzik:2024wtq and QCD phase transition Byrnes:2018clq . The main force to resist the collapse of an overdense region comes from the gas pressure, which is reduced with the decrease of the equation of state during a phase transition. Therefore, the probability of PBH formation increases with the decrease of $\omega$. Eq. (6) shows that domain walls with $\overline{v}_{w}\simeq 0.4$ in the scaling regime have a negative equation of state, $\omega\simeq-2/3$. This is because domain walls have a negative surface pressure for $\overline{v}_{w}\lesssim 0.8$ (corresponding to $\overline{v}_{w}^{2}\gamma^{2}\simeq 2$), as shown by Eq. (4). We define an effective equation of state for the mixture gas of radiation and domain walls:
$\omega_{\rm eff}=\frac{p_{w}+p_{\rm rad}}{\rho_{w}+\rho_{\rm rad}}=\frac{1-2f_{w}}{3(1+f_{w})}.$ (48)
The cosmic equation of state can get smaller when the domain walls become a nontrivial component of the universe. The variation of the critical density contrast $\delta_{c}$ as a function of the equation of state has been calculated in Ref. Musco:2012au . We find these simulation results can be reproduced by the following fit function
$\delta_{c}=\frac{A_{0}}{\left[1+(d_{0}/\omega_{\rm eff})^{s_{1}}\right]^{s_{2}}},$ (49)
where the fit parameters are provided in Tab. 1. In Fig. 3, we present the fit results of the simulations with different values of the perturbation profile parameter $\alpha$. We observe that all the simulation results are nicely reproduced with the fit formula. In this work, we adopt $\alpha=0$ as the default choice, which represents a Gaussian curvature profile.
Figure 3: The solid curves represent the fit to the simulation results, denoted by the colored points, given in Ref. Musco:2012au with Eq. (49).
Parameter $A_{0}$ $d_{0}$ $s_{1}$ $s_{2}$
$\alpha=0$ 0.65188 0.11922 0.94511 1.12954
$\alpha=1$ 0.67491 0.10604 0.93996 1.16995
$\alpha=2$ 0.67756 0.1109 0.97455 1.09519
$\alpha=5$ 0.69254 0.10163 0.94824 1.15388
Table 1: Fit results of simulations in Ref. Musco:2012au with the fit formula (49) for different values of $\alpha$.
Figure 4: Left: The critical density contrast $\delta_{c}$ as a function of the PBH mass $M$. Right: The PBH relic abundance $\Omega_{\rm PBH}$ as a function of the PBH mass $M$, assuming $N=1$ and a constant variance $\sigma^{2}=0.004$. The black lines represent the results with a constant $\delta_{c}$ in the radiation phase. The effects on $\delta_{c}$ and $\Omega_{\rm PBH}$ from the domain walls are represented by the blue and green curves for $\sigma_{w}^{1/3}=10^{5}$ GeV and $10^{6}$ GeV, respectively.
In the left plot of Fig. 4, we show the critical density contrast $\delta_{c}$ in Eq. (49) as a function of the PBH mass. We observe that $\delta_{c}$ sharply decreases when $M\sim 10^{1}\leavevmode\nobreak\ M_{\odot}$ (blue curve with $\sigma_{w}^{1/3}=10^{5}$ GeV) and $M\sim 10^{-2}\leavevmode\nobreak\ M_{\odot}$ (green curve with $\sigma_{w}^{1/3}=10^{6}$ GeV), corresponding to $f_{w}\sim 10^{-1}$. This leads to an enormous improvement in the PBH formation probability, as shown in the right plot of Fig. 4, where we adopt a constant variance $\sigma^{2}=0.004$. Therefore, once a fraction $f_{w}\sim 10^{-1}$ of the cosmic energy density is composed of the domain walls, the softened equation of state by the domain wall could also lead to a significant enhancement of the PBH relic abundance.
7 Gravitational waves from domain wall annihilation
It is known that the domain wall’s energy density scales as $\propto t^{-1}$ while the energy density of radiation scales as $t^{-2}$. As a result, the domain walls would dominate the evolution of the universe if they existed for long enough. A common solution to the problem is to introduce a bias potential which lifts the degenerate vacua by an energy $V_{\rm bias}=V_{\rm flase}-V_{\rm true}$ so that the domain walls can annihilate at some time $t_{\rm ann}$ before it over-close the universe. The annihilation of the domain wall takes place when the bias energy becomes comparable to the surface energy of the domain wall, i.e., $V_{\rm bias}\simeq\rho_{w}$, which gives the annihilation temperature
$T_{\rm ann}\simeq 334.7\leavevmode\nobreak\ {\rm MeV}\left(\frac{1}{\mathcal{A}}\right)^{1/2}\left(\frac{10.75}{g_{*}(T_{\rm ann})}\right)^{1/4}\left(\frac{V_{\rm bias}^{1/4}}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)^{2}\left(\frac{10^{5}\leavevmode\nobreak\ {\rm GeV}}{\sigma_{w}^{1/3}}\right)^{3/2}.$ (50)
This corresponds to a horizon mass
$M_{H}(T_{\rm ann})\simeq\frac{\mathcal{A}\sigma_{w}}{GV_{\rm bias}}=1.34\leavevmode\nobreak\ M_{\odot}\left(\frac{\mathcal{A}}{1}\right)\left(\frac{0.1\leavevmode\nobreak\ {\rm GeV}}{V_{\rm bias}^{1/4}}\right)^{4}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ {\rm GeV}}\right)^{3}.$ (51)
Simulations have indicated that the emissions of gravitational waves and radiation from domain walls in the scaling regime are negligible. Nearly all the domain wall energy is converted into gravitational waves when the domain walls are destroyed by the bias potential. The peak frequency of gravitational waves is obtained at the annihilation time of the domain walls, given by $f_{\rm peak}(t_{\rm ann})\simeq H(t_{\rm ann})$, and is red-shifted by the cosmic expansion to the present as
$f_{\rm peak}(t_{0})\simeq 1.21\times 10^{-8}\mathrm{\leavevmode\nobreak\ Hz}\left(\frac{g_{*}\left(T_{\mathrm{ann}}\right)}{10.75}\right)^{1/2}\left(\frac{g_{*s}\left(T_{\mathrm{ann}}\right)}{10.75}\right)^{-1/3}\left(\frac{T_{\mathrm{ann}}}{0.1\leavevmode\nobreak\ \rm{GeV}}\right).$ (52)
The peak gravitational wave amplitude produced at the annihilation time is determined as
$\Omega_{\mathrm{GW}}\left(f_{\mathrm{peak}}\left(t_{\mathrm{ann}}\right)\right)\simeq\frac{8\pi\tilde{\epsilon}_{\mathrm{gw}}G^{2}\mathcal{A}^{2}\sigma_{w}^{2}}{3H^{2}(t_{\rm ann})},$ (53)
with $\tilde{\epsilon}_{\rm{gw}}\simeq 0.7\pm 0.4$ Hiramatsu:2013qaa . The peak gravitational wave amplitude today is estimated as
$h^{2}\Omega_{\rm{GW}}^{\rm peak}\left(t_{0}\right)=2.4\times 10^{-10}\left(\frac{\tilde{\epsilon}_{\rm{gw}}}{0.7}\right)\left(\frac{\mathcal{A}}{1}\right)^{2}\left(\frac{g_{*,s}(T_{\rm ann})}{10.75}\right)^{-4/3}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ {\rm GeV}}\right)^{6}\left(\frac{T_{\rm ann}}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)^{-4}.$ (54)
Using Eq. (45) we have
$\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ {\rm GeV}}\right)^{3}=39.6f_{w}(T)\left(\frac{1}{\mathcal{A}}\right)\left(\frac{g_{*}(T)}{10.75}\right)^{1/2}\left(\frac{T}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)^{2}.$ (55)
Combining Eqs. (54) and (55), we have
$h^{2}\Omega_{\rm{GW}}^{\rm peak}\left(t_{0}\right)=3.8\times 10^{-9}\left(\frac{\tilde{\epsilon}_{\rm{gw}}}{0.7}\right)\left(\frac{g_{*,s}(T_{\rm ann})}{10.75}\right)^{-4/3}\left(\frac{g_{*}(T_{\rm ann})}{10.75}\right)\left(\frac{f_{w}(T_{\rm ann})}{0.1}\right)^{2}.$ (56)
The causality requires the gravitational wave spectrum scale as $\propto(f/f_{\rm peak})^{3}$ for the frequency below $f_{\rm peak}$ Caprini:2009fx . On the other hand, for $f\gtrsim f_{\rm peak}$, numerical simulations Hiramatsu:2013qaa indicate a scaling behavior $\propto(f/f_{\rm peak})^{-1}$. We adopt the following fit formula for the gravitational wave spectrum from the domain wall annihilation Ferreira:2022zzo
$h^{2}\Omega_{\rm{GW}}=h^{2}\Omega_{\rm{GW}}^{\rm peak}\frac{(\gamma_{c}+c_{1})^{c_{2}}}{(c_{1}x^{-\gamma_{c}/c_{2}}+\gamma_{c}x^{c_{1}/c_{2}})^{c_{2}}},$ (57)
where $x=f/f_{\rm peak}$, $\gamma_{c}=3$ from the requirement of causality, and $c_{1},\leavevmode\nobreak\ c_{2}\simeq 1$ from the simulations. However, $c_{1}$ and $c_{2}$ may depend on the physics model, as indicated by the simulations Hiramatsu:2012sc .
8 PTA data analysis
The recent PTA observations from NANOGrav NANOGrav:2023gor ; NANOGrav:2023hvm , CPTA Xu:2023wog , EPTA EPTA:2023fyk , and PPTA Reardon:2023gzh provide $\sim 3\sigma$ evidence for the existence of stochastic gravitational wave background (SGWB) at the nano-Hz frequency band. Various scenarios have been proposed to interpret the SGWB signal (see e.g., Madge:2023dxc ; Bian:2023dnv ; Wu:2023hsa ; Ellis:2023oxs for a summary). The domain wall annihilation is a compelling interpretation for PTAs because the bias potential induced by the QCD phase transition can naturally lead to nano-Hz gravitational waves from the annihilation of domain walls Chiang:2020aui ; Lu:2023mcz . It has been noticed in Ref. Lu:2023mcz that the domain wall should annihilate before it dominates the Universe so that the gravitational wave signal could be loud enough for the PTA experiments. As shown in this work, this would also lead to significant PBH productions due to the domain wall number density fluctuations and the equation of state reduction. Our PBH production mechanism is different from those by the collapse of closed domain walls, which depends heavily on the annihilation process of the walls Ipser:1983db ; Ferrer:2018uiu .
For our purpose, we follow the standard Bayesian statistics to analyze the PTA data, including the IPTA-DR2 dataset IPTA:datalink and the recent NG15 dataset NANOGrav:2023gor . The timing-residual cross-power spectral density measured by the PTAs is given by
$S_{ab}(f)=\Gamma_{ab}\frac{h_{c}^{2}(f)}{12\pi^{2}f^{3}},$ (58)
where the characteristic strain $h_{c}(f)$ is related to the gravitational wave background relic abundance via
$\Omega_{\mathrm{GW}}(f)=\frac{2\pi^{2}f^{2}h_{c}^{2}(f)}{3H_{0}^{2}}.$ (59)
We consider an isotropic and unpolarized GW background (GWB) overlap reduction function $\Gamma_{ab}$, which is described by the Hellings-Downs function.
We employ the package PTArcade Mitridate:2023oar , which wraps the code enterprise Ellis:2020zenodo and ceffyl Lamb:2023jls , to implement the GW signal model. The Bayesian inference is adopted here to derive information on the parameters by fitting to the pulsars’ timing residuals $\delta t$, which is given by the sum of white noise and red noise, including the pulsar-intrinsic red noise and common red noise produced by a GWB (see appendix for more detail). Furthermore, the package PTMCMCSampler Ellis:2017 is employed to perform the Markov chain Monte Carlo (MCMC) sampling. In the fit to each dataset, we separately generate $2\times 10^{6}$ MC sample points. The parameter posterior distributions are derived using GetDist Lewis:2019xzd . Following Refs. Antoniadis:2022pcn ; NANOGrav:2023gor to avoid the pulsar-intrinsic excess noise at high frequencies, only the first 13 and 14 frequencies of the IPTA-DR2 and NG15 datasets are included in our analysis. The priors of the parameters for signal model and noise are provided in table 3, together with the posteriors from the analysis presented in table 4 and Fig. 11 in appendix E. It has been previously found that the domain wall scenario can improve the fit to the NANOGrav data by a Bayesian factor $B\simeq 10$ NANOGrav:2023hvm . The purpose of our fit in this work is not to repeat the model comparison but to estimate the parameters regions of our concerns.
Figure 5: The red (light red) and blue (light blue) regions represent the $1\sigma$ ($2\sigma$) regions of the 2D posterior distributions of the annihilation temperature $T_{\rm ann}$ and the domain wall density fraction $f_{w}(\rm ann)$ by fit to the NG15 and IPTA-DR2 datasets, respectively. The black dashed lines represent the $\Delta N_{\rm eff}$ constraints. The solid green line and dotted gray line represent contours of $\Omega_{\rm PBH}=10^{-3}$ and $0.25$ from the domain wall fluctuations, respectively.
Figure 6: The same as Fig. 5, with $N=6$ (left panel) and $N=10$ (right panel).
In Fig. 5 and Fig. 6, we show the interpretations of the domain wall scenario for the PTA observation, noting that our fit based on Eqs (52) and (56) is $N$-independent. The red and blue regions represent the 2D posterior distributions of $T_{\rm ann}$ and $f_{w}(T_{\rm ann})$ by fit to the NG15 and IPTA-DR2 datasets, respectively. In addition, the fit to the observations indicates that a significant cosmic energy density consists of the domain walls. The relativistic quanta released by the domain walls before their annihilation could leave an observable imprint on the standard cosmology. In order not to spoil the observations of BBN and CMB, we require the following conditions on the domain wall scenario, depending on the specific model:
  • Case I: Domain walls decay to standard model particles. In this case, the domain wall should annihilate before the onset of BBN at $T_{\rm BBN}\simeq 2.7$ MeV.
  • Case II: Domain walls mainly decay to dark radiation in a dark sector, which weakly contacts with or totally decouples from the visible plasma. The dark radiation could alter the expansion rate of the universe by contributing to the effective number of neutrino species $\Delta N_{\rm eff}=\rho_{\rm DR}(t_{\rm rec})/\rho_{\nu}(t_{\rm rec})$, where $t_{\rm rec}$ is the epoch of recombination and the energy density of a single neutrino species is given by
    $\rho_{\nu}(t_{\rm rec})=2\times\frac{7}{8}\times\frac{\pi^{2}}{30}\left(\frac{4}{11}\right)^{4/3}T_{\rm rec}^{4},$ (60)
    where the recombination temperature $T_{\rm rec}\simeq 0.3$ eV. With Eqs. (22), (33), and (39) the effective number can be estimated as
    $\Delta N_{\rm eff}$ $=$ $\frac{\rho_{\rm DR}(t_{\rm ann})}{\rho_{\nu}(t_{\rm rec})}\left(\frac{g_{*,s}(T_{\rm rec})T_{\rm rec}^{3}}{g_{*,s}(T_{\rm ann})T_{\rm ann}^{3}}\right)^{4/3}$ (61)
    $\simeq$ $f_{w}(T_{\rm ann})\left(\frac{p_{\rm DR}}{0.16}\right)\left(\frac{g_{*}(T_{\rm ann})}{10.75}\right)\left(\frac{10.75}{g_{*,s}(T_{\rm ann})}\right)^{4/3},$
    where we have used $g_{*,s}(T_{\rm rec})\simeq 3.91$ and $\rho_{\rm DR}(T_{\rm ann})=p_{\rm DR}f_{w}\rho_{\rm rad}(T_{\rm ann})$, with $0\leq p_{\rm DR}\leq 1$ being the energy fraction of domain walls that is transferred to dark radiation. We adopt $p_{\rm DR}=1$ for a conservative estimate.
For case I, Fig. 5 and Fig. 6 show that to interpret PTAs, the annihilation of domain walls should take place during the QCD scale before the onset of BBN, i.e., the annihilation temperature $T_{\rm ann}$ should fall in the range of $\sim 0.01-1$ GeV. Therefore, domain walls do not affect the BBN processes if the domain wall decays mainly into standard model particles. For case II, $\Delta N_{\rm eff}$ observations can put useful constraints if the domain wall decays into dark radiation. The current CMB from Planck and BAO observations provide an upper bound on the number of extra neutrino species $\Delta N_{\rm eff}\leq 0.29$ at 95% confidence level (CL) Planck:2018vyg (the black dashed lines). The upcoming CMB experiments including Simons Observatory SimonsObservatory:2018koc and the CMB-S4 experiment CMB-S4:2022ght could further tighten this upper bound to 0.11 and 0.06, respectively. Indeed, numerical estimates have indicated that at most about $(10-20)\%$ of the wall energy $\rho_{w}$ would radiate as relativistic quanta during its evolution in the scaling regime. Therefore, constraints from extra neutrino species by Eq. (61) with $p_{\rm DR}=1$ are less conservative. We expect the observations from future Simons Observatory and CMB-S4 experiments can probe part of the parameter space in case II.
[…]
In Fig. 7, we show the fit results for the IPTA-DR2 and NG15 gravitational spectra. We observe that the domain wall scenario can well explain the PTA observations while a large uncertainty remains at frequencies $f\gtrsim 10$ nHz. Note that the domain walls annihilate to gravitational radiation before BBN can also make a contribution to the extra neutrino species and affect the expansion rate of the universe. The bound on the relic abundance of gravitational radiation today is related to the effective number of neutrino species by Caprini:2018mtu
$h^{2}\Omega_{\mathrm{GW}}(t_{0})\lesssim 5.6\times 10^{-6}\Delta N_{\mathrm{eff}}.$ (62)
The gray region and red dashed line represent the upper bounds of the gravitational radiation relic abundance from Planck+BAO and future CMB-S4, respectively. The gravitational spectra at frequencies $\gtrsim 10^{-7.5}-10^{-7}$ Hz might have some overlaps with the CMB-S4 experiment, although the pulsar-intrinsic excess noise exists in this frequency band. Therefore, the upcoming CMB experiments will shed more light on the nano-Hz SGWB.
Combining Eq. (50) and Eq. (55), we obtain the bias potential in terms of $f_{w}(T_{\rm ann})$ and $T_{\rm ann}$
$\frac{V_{\rm bias}^{1/4}}{0.1\leavevmode\nobreak\ {\rm GeV}}=1.37(f_{w}(T_{\rm ann}))^{1/4}\left(\frac{T_{\rm ann}}{0.1\leavevmode\nobreak\ {\rm GeV}}\right)\left(\frac{g_{*}(T_{\rm ann})}{10.75}\right)^{1/4}.$ (63)
Together with Eq. (55) we can therefore transfer the fit results in the $f_{w}-T_{\rm ann}$ space to the $\sigma_{w}-V_{\rm bias}$ plane. We show these results in Fig. 8, which indicate that the typical bias potential $V_{\rm bias}^{1/4}\sim\Lambda_{\rm QCD}\sim 0.1$ GeV and the surface tension $\sigma_{w}^{1/3}\sim 10^{5}$ GeV and they present a strong correlation.
Figure 8: $1\sigma$ and $2\sigma$ contours of posterior distributions in the $\sigma_{w}^{1/3}-V_{\rm bias}^{1/4}$ plane, with $N=1$ and $N=10$ in the left and right panels, respectively.
9 Constraints and implications from PBH formation
Most of the domain wall energy $\rho_{w}$ is released at the annihilation temperature $T_{\rm ann}$, beyond which $\rho_{w}$ scales as $T^{\alpha}$ as indicated by the simulations. Let us parametrize the temperature-dependent domain wall energy density as follows Kawasaki:2014sqa ; Ferrer:2018uiu
$\rho_{w}(T)=\left\{\begin{matrix}\rho_{w}&\leavevmode\nobreak\ \leavevmode\nobreak\ {T\geq T_{\rm ann}},\\ \rho_{w}\left(\frac{T}{T_{\rm ann}}\right)^{\alpha}&\leavevmode\nobreak\ \leavevmode\nobreak\ {T<T_{\rm ann}},\end{matrix}\right.$ (64)
where $\rho_{w}=\mathcal{A}\sigma_{w}/t$. A fit to the simulation of the domain wall evolution gives $\alpha\simeq 7$ Kawasaki:2014sqa . Then the evolution of the fraction of domain wall energy density is given by
$f_{w}=\left\{\begin{matrix}1.81\times 10^{-3}\frac{\mathcal{A}}{1}\frac{M_{H}(T)}{M_{\odot}}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ \rm GeV}\right)^{3}&M_{H}\leq M_{H}(T_{\rm ann})\\ 1.81\times 10^{-3}\frac{\mathcal{A}}{1}\frac{M_{H}(T)}{M_{\odot}}\left(\frac{\sigma_{w}^{1/3}}{10^{5}\leavevmode\nobreak\ \rm GeV}\right)^{3}\left(\frac{M_{H}(T_{\rm ann})\sqrt{g_{*}(T_{\rm ann})}}{M_{H}(T)\sqrt{g_{*}(T)}}\right)^{\alpha/2}&M_{H}>M_{H}(T_{\rm ann}).\end{matrix}\right.$ (65)
It is worth mentioning that, unlike the PBHs from the collapse of closed domain walls, our mechanism is not sensitive to the evolution of the network after the onset of domain wall annihilation, whose details remain to have a large uncertainty Ferrer:2018uiu .
In Fig. 9, we show the PBH relic abundance $\Omega_{\rm PBH}(f_{w},T)$ from the domain wall number density fluctuations, with the evolution of wall energy fraction according to Eq. (65). The annihilation temperatures $T_{\rm ann}$ and the energy fractions $f_{w}$ at $T_{\rm ann}$ are taken to be (0.1GeV, $(0.1\leavevmode\nobreak\ {\rm GeV},\leavevmode\nobreak\ 0.1)$ (blue curves) and (0.5GeV, $(0.5\leavevmode\nobreak\ {\rm GeV},\leavevmode\nobreak\ 0.2)$ (green curves), which are two benchmarks for PTA observations. The solid and dashed curves correspond to the results with the domain wall number $N=1$ and $N=4$, respectively. The surface tension can be determined by Eq. (55) at $T_{\rm ann}$ and the horizon mass at the annihilation temperature is $M_{H}(T_{\rm ann})=1/(2GH(T_{\rm ann}))$. For our choices, the PBH masses at annihilation are $M(0.1\leavevmode\nobreak\ {\rm GeV})\simeq 0.12M_{\odot}$ (orange vertical line) and $M(0.5\leavevmode\nobreak\ {\rm GeV})\simeq 3.0M_{\odot}$ (red vertical line). We observe that the PBH relic abundance $\Omega_{\rm PBH}$ has a narrow spike around $M(T_{\rm ann})$, below and above which $\Omega_{\rm PBH}$ sharply decreases with the PBH mass. Therefore, PBHs are significantly produced at the moment close to domain wall annihilation. This fact will be adopted to put useful constraints on the model parameter space of the domain wall scenario.
[…]
Parameter $N=1$ $N=4$ $N=6$ $N=10$
$f_{w}(T_{\rm ann}=0.1\leavevmode\nobreak\ {\rm GeV})$ 0.07 0.12 0.15 0.18
Table 2: Upper limits on the domain wall energy fraction $f_{w}$ from the PBH abundance, with various values of the domain wall number $N$.
From Eq. (40) and Fig. 10 we observe that to satisfy various experimental constraints on the PBH relic abundance for $M\sim M_{\odot}$, we should require $\beta(M)\lesssim 10^{-8}-10^{-7}$, corresponding to $x=\delta_{c}/(\sqrt{2}\sigma)\gtrsim 4$. The density fluctuations deep in the tails of the probability distribution exponentially suppress the production of PBHs via the critical density collapse mechanism (see Fig. 10 in appendix D). The parameters that give the probability of a density contrast within $3\sigma$ are strongly constrained by the PBH abundance. To estimate these bounds, we note that the PBH abundance has a sharp spike at the annihilation temperature. Therefore, we calculate the $f_{w}$-dependent critical density contrast $\delta_{c}$ with Eqs. (48) and (49) and determine $\Omega_{\rm PBH}(f_{w}(T_{\rm ann}),T_{\rm ann})$ with Eq. (40) at $T_{\rm ann}$. In table 2, we provide the upper limits on the domain wall energy fraction $f_{w}$ at the annihilation temperature $T_{\rm ann}=0.1$ GeV from the condition $\Omega_{\rm PBH}\leq 10^{-3}$ for various values of the domain wall number $N$. In Fig. 5 and Fig. 6, we show the contours of PBH relic abundance with $\Omega_{\rm PBH}=10^{-3}$ (green solid line) and $\Omega_{\rm PBH}=0.25=\Omega_{\rm DM}$ (gray dot line) in the $f_{w}(T_{\rm ann})-T_{\rm ann}$ plane. We observe from Fig. 5 that the scenario for PTAs with the domain wall number $N=1$ (corresponding to models with a $Z_{2}$ symmetry) has been excluded by the PBH abundance from domain wall number density fluctuations. A marginal space out of $2\sigma$ CL remains available for the case with $N=2$. We therefore also exclude such models for explaining the nano-Hz SGWB. For the case of $N=3$ and $N=4$, most of the parameter space for NG15 and about half of the space region for IPTA-DR2 are excluded by the PBH relic abundance. Finally, for models with $N=6$ and $N=10$, we show in Fig. 6 that most of the parameter space for IPTA-DR2 and about half of the space region for NG15 are not constrained by the PBH formation. Models for $N\sim 10$ domain walls include the clockwork axion, in which several domain walls are produced following the spontaneous breakdown of a set of discrete shift symmetries.
10 Conclusions and discussions
In this work, we propose that PBHs can be formed from the number density fluctuations of the domain walls in a late stage. The network consisting of a horizon-sized domain wall with a small number density could lead to large Poisson fluctuations. Since the energy density of the domain wall scales as $t^{-1}$ in the scaling regime, the probability for the production of horizon-size density perturbations with large amplitudes from domain wall number density fluctuations grow with time. Once the density contrast exceeds the critical value, it stops expanding and collapses to form PBHs. The outburst of PBH formation takes place when the fraction of domain wall energy density increases to $f_{w}\sim 0.1$. In addition, the domain wall with negative pressure could soften the equation of state and reduce the gas pressure that resists the collapse of density perturbations and, therefore, increase the probability of PBH formation.
By a fit to the NANOGrav and IPTA-DR2 datasets, we show that the gravitational waves from domain wall annihilation could well explain the nano-Hz SGWB if the domain walls account for $\sim 10$% of the total energy density at the annihilation temperature $T_{\rm ann}\sim 0.1$ GeV. Thus, the formation of PBHs from domain wall number density fluctuations before the annihilations can be significant. We find that the overproduction of PBH abundance from the fluctuations of the network with the domain wall numbers $N=1$ and $N=2$ have constrained the parameter space for the interpretation of PTA observations. We therefore exclude the domain wall scenario for the PTAs in models with a $Z_{2}$ symmetry. Some of the parameter space for NG15 is constrained in models with $N\lesssim 10$. With these results, we conclude that scenarios for NG15 in terms of an explicit $Z_{\mathcal{N}}$ symmetry King:2023ayw ; Pham:2024vso in a model with a soft $U(1)$ symmetry breaking term Chiang:2019oms ; Bai:2023cqj
$V(\Phi)\supset\frac{1}{M^{\mathcal{N}-4}}(\Phi^{\mathcal{N}}+\Phi^{\dagger\mathcal{N}}),$ (66)
where $\Phi$ is a complex scalar $\Phi$, $M$ has a mass dimension, and $\mathcal{N}\leq 4$ from the renormalizability, are strongly constrained.
To realize a model with $\mathcal{N}\sim 10$ domains (and the domain wall number $N\sim\mathcal{N}$), one can consider $\mathcal{N}$ copies of complex scalars $\Phi_{i}$ (with $i=1,...,\mathcal{N}$). The global $U(1)^{\mathcal{N}}$ symmetry is spontaneously broken when each of $\Phi_{i}$ obtains a common vacuum expectation value. We can further explicitly break the $\mathcal{N}$ $U(1)$ symmetries to $\mathcal{N}-1$ discrete shift symmetries and one $U(1)$ symmetry by a term $\Phi_{i}^{\dagger}\Phi_{i+1}^{q}$, which then leads to the formation of $\mathcal{N}-1$ domains. Such symmetries have been naturally embedded in the framework of the clockwork axion models Kaplan:2015fuy and deconstruction of extradimensional grand unified theories (GUTs) Arkani-Hamed:2001kyx ; Li:2002xd .
Another realization of the $Z_{\mathcal{N}}$ symmetry is the axion-like models (see, for example, Ref. ZambujalFerreira:2021cte ). Consider a hidden non-Abelian $SU(n)$ gauge sector consisting of a complex scalar $\Phi$ and a set of vector-like quarks. The potential of $\Phi$ possesses a global $U(1)$ symmetry, which is spontaneously broken by the Higgs mechanism. If the vector-like quarks are charged under the $U(1)$ transformation, then the quantum anomaly of $U(1)$ with respect to the $SU(n)$ gauge group can break the $U(1)$ symmetry to its $Z_{\mathcal{N}}$ subgroup via the instanton effect during the confinement of the $SU(n)$ symmetry, which gives rise to an axion-like periodic potential
$V(a)\simeq\Lambda_{n}^{4}\left(1-\cos\left(\frac{\mathcal{N}a}{v}\right)\right),$ (67)
where $a$ denotes the axion-like particle, $\Lambda_{n}\gg\Lambda_{\rm QCD}$ is the $SU(n)$ confinement scale, and $v$ is the vacuum expectation value of the scalar.
We also note that our constraints can be applied to models with the domain wall number $N\lesssim 10$. For models with the domain wall number $N\gg 10$, the production of PBHs from the domain wall number density fluctuations is strongly suppressed by the Poisson statistics and, therefore, the constraints are insignificant.
Finally, let’s briefly review uncertainties that commonly exist in the critical collapse PBH formation mechanism Carr:2016drx . (i) Non-sphericity: We have assumed a homogeneous and spherically symmetric perturbation. The non-sphericity of the perturbation can improve the threshold and reduce the possibility for PBH formation Kuhnel:2016exn . In our scenario, the non-sphericality of the overdense region induced by the domain wall fluctuations is expected to be small since the radiation always dominates the universe. (ii) Non-Gaussianity: We have assumed a Gaussian probability density distribution for the density contrast. The critical collapse mechanism for PBH formation relies on the exponential suppression in the upper tail of the probability distribution. Therefore, a non-Gaussian component in the distribution could exponentially change the abundance of PBH. Further investigations on the non-Gaussianity in the domain wall fluctuations are needed. Nevertheless, the fluctuations do not affect the CMB-scale observations if the domain walls annihilate long before the onset of BBN.
11 Acknowledgments
BQL is supported in part by Zhejiang Provincial Natural Science Foundation of China under Grant No. LQ23A050002. CWC is supported in part by the National Science and Technology Council under Grant No. NSTC-111-2112-M-002-018-MY3. TL is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201504, by the Project No. 12275333 supported by the National Natural Science Foundation of China, by the Scientific Instrument Developing Project of the Chinese Academy of Sciences, Grant No. YJKYYQ20190049, and by the International Partnership Program of Chinese Academy of Sciences for Grand Challenges, Grant No. 112311KYSB20210012.
Appendix A Domain wall energy tension in toy model
Let’s consider the following toy model for the real scalar $\phi$ as an example Vilenkin:1984ib ; Gelmini:1988sf ; Saikawa:2017hiv :
$\mathcal{L}=\frac{1}{2}\left(\partial_{\mu}\phi\right)^{2}-V(\phi),$ (68)
where
$V(\phi)=\frac{1}{4}\lambda\left(\phi^{2}-v^{2}\right)^{2}.$ (69)
The $Z_{2}$ symmetry of the model is spontaneously broken when $\phi$ obtains vacuum expectation values ±$\pm v$. The manifold $\mathcal{M}$ of the degenerate vacua contains two points, $\langle\phi\rangle=+v$ and $\langle\phi\rangle=-v$, which are separated by the domain wall of the false vacuum with $\langle\phi\rangle=0$.
Using the Lagrangian density (68), the description of the domain wall can be obtained by solving the equation of motion
$\square\phi+\lambda\phi\left(\phi^{2}-v^{2}\right)=0,$ (70)
where the d’Alambertian $\square=\partial_{\mu}\partial^{\mu}=\partial_{0}^{2}-\nabla^{2}$. Suppose the domain wall is infinite and lies in the $yz$ plane at $x=0$. The solution that satisfies the boundary conditions $\phi(x=+\infty)=+v$ and $\phi(x=-\infty)=-v$ is found to be
$\phi(z)=v\tanh(x/\delta),$ (71)
where $\delta=1/(\sqrt{\lambda/2}v)$ denotes the thickness of the wall. The energy-momentum tensor for the domain wall solution (71) is given by
$T_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi-g_{\mu\nu}\mathcal{L}.$ (72)
With the wall configuration (71) and $\partial_{\mu}\phi(x)=0$ for $\mu\neq x$, we have
$T_{00}=-T_{22}=-T_{33}=\frac{1}{2}\left(\partial_{x}\phi\right)^{2}+V(\phi)\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ T_{11}=0.$ (73)
Then we obtain
$T_{\mu\nu}=\frac{\lambda}{2}v^{4}[{\rm sech}(x/\delta)]^{4}\operatorname{diag}(1,0,-1,-1).$ (74)
We can define the energy-momentum surface density tensor of the walls as
[…]
Using Eq. (74), the surface energy density of the wall is given by
$\sigma_{w}=S_{00}=-S_{22}=-S_{33}=\frac{4}{3}\sqrt{\frac{\lambda}{2}}v^{3},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ S_{11}=0.$ (76)
Appendix B Surface energy-momentum tensor
In the wall rest frame $K^{\prime}$, the surface tensor is given by Eqs. (75) and (76). In the observer’s rest frame $K$, $K^{\prime}$ moves along the $\mathbf{x}$ axis with velocity $\overline{v}_{w}$, then the surface density tensor $S_{\mu\nu}^{x}(\overline{v}_{w})$ measured by the observer is obtained by performing the Lorentz transformation
$S_{00}^{x}$ $=\gamma^{2}(S_{00}^{\prime}+2\overline{v}_{w}S_{01}^{\prime}+{\overline{v}_{w}}^{2}S_{11}^{\prime}),$ (77)
$S_{11}^{x}$ $=\gamma^{2}(S_{11}^{\prime}+2\overline{v}_{w}S_{01}^{\prime}+{\overline{v}_{w}}^{2}S_{00}^{\prime}),$
$S_{01}^{x}$ $=S_{10}^{x}=\gamma^{2}(S_{01}^{\prime}(1+{\overline{v}_{w}}^{2})+\overline{v}_{w}S_{00}^{\prime}+{\overline{v}_{w}}S_{11}^{\prime}),$
$S_{22}^{x}$ $=S_{22}^{\prime},\leavevmode\nobreak\ \leavevmode\nobreak\ S_{33}^{x}=S_{33}^{\prime},$
where $\gamma\equiv 1/\sqrt{1-\overline{v}_{w}^{2}}$ and denotes the value in the $K^{\prime}$ frame. The wall plane may also move along the $\mathbf{y}$ or $\mathbf{z}$ axis. The surface energy-momentum tensors measured by the observer are
$S_{\mu\nu}^{x}=\left(\begin{array}[]{cccc}\gamma^{2}\sigma_{w}&\gamma^{2}\sigma_{w}\overline{v}_{w}&0&0\\ \gamma^{2}\sigma_{w}\overline{v}_{w}&\gamma^{2}\sigma_{w}\overline{v}_{w}^{2}&0&0\\ 0&0&-\sigma_{w}&0\\ 0&0&0&-\sigma_{w}\end{array}\right),$ (78)
$S_{\mu\nu}^{y}=\left(\begin{array}[]{cccc}\gamma^{2}\sigma_{w}&0&\gamma^{2}\sigma_{w}\overline{v}_{w}&0\\ 0&-\sigma_{w}&0&0\\ \gamma^{2}\sigma_{w}\overline{v}_{w}&0&\gamma^{2}\sigma_{w}\overline{v}_{w}^{2}&0\\ 0&0&0&-\sigma_{w}\end{array}\right),$ (79)
$S_{\mu\nu}^{z}=\left(\begin{array}[]{cccc}\gamma^{2}\sigma_{w}&0&0&\gamma^{2}\sigma_{w}\overline{v}_{w}\\ 0&-\sigma_{w}&0&0\\ 0&0&-\sigma_{w}&0\\ \gamma^{2}\sigma_{w}\overline{v}_{w}&0&0&\gamma^{2}\sigma_{w}\overline{v}_{w}^{2}\end{array}\right).$ (80)
In addition, in each axis, there are two directions available for the wall plane. The corresponding tensors can be obtained by the replacement $\overline{v}_{w}\to-\overline{v}_{w}$ in the above results. Averaged over the six directions, the off-diagonal terms cancel out, and the average surface energy tensor
$S_{\mu\nu}(\overline{v}_{w})=\frac{\sigma_{w}}{3}{\rm diag}(\gamma^{2},\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2,\overline{v}_{w}^{2}\gamma^{2}-2).$ (81)
Appendix C Wall evolution model
We follow Ref. Martins:2016book to review the VOS model for describing the evolution of the wall. Let us start from the world volume (Dirac) action Martins:2016book
$S=-\int\mathcal{L}d^{3}\sigma=-\sigma_{w}\int\sqrt{\gamma}d^{3}\sigma,$ (82)
where $\gamma=\frac{1}{3!}\varepsilon^{ab}\varepsilon^{cd}\gamma_{ac}\gamma_{bd}$, with $\gamma_{ab}=g_{\mu\nu}x_{,a}^{\mu}x_{,b}^{\nu}$ and $x_{,a}^{\mu}=\frac{\partial x^{\mu}}{\partial\sigma^{a}}$, $x_{\mu}$ and $\sigma_{a}$ are respectively the spacetime coordinates and worldsheet (two-dimensional surface) coordinates, $g_{\mu\nu}$ and $\gamma_{ab}$ are respectively the 4D spacetime and 2D worldsheet metrics. The equation of motion (EOM) in the flat FRW universe is given by
$\frac{\dot{a}}{a}\delta_{0\lambda}\sqrt{\gamma}\gamma^{ab}\gamma_{ab}-\partial_{c}\left(\sqrt{\gamma}\gamma^{ab}g_{\mu\lambda}x_{,a}^{\mu}\delta_{b}^{c}\right)=0.$ (83)
Introduce the averaged energy density and the root-mean-squared (RMS) velocity of the wall
$\bar{\rho}_{w}\equiv\frac{E}{V}=\frac{\sigma_{w}a^{2}}{V}\int\varepsilon d^{2}\sigma,\quad\overline{v}_{w}^{2}\simeq v_{\rm rms}^{2}=\frac{\int\dot{x}^{2}\varepsilon d^{2}\sigma}{\int\varepsilon d^{2}\sigma},$ (84)
where $\varepsilon=\sqrt{\gamma}\gamma^{00}/a$ with $a$ being the cosmic scale factor. By averaging the temporal and spatial components of the EOM (83), we obtan Martins:2016book
$\frac{d\overline{\rho}_{w}}{dt}$ $=-H\overline{\rho}_{w}\left(1+3\overline{v}_{w}^{2}\right)=-3(1+w)H\overline{\rho}_{w},$ (85)
$\frac{d\overline{v}_{w}}{dt}$ $=\left(1-\overline{v}_{w}^{2}\right)\left(\frac{k_{w}}{\overline{L}}-3H\overline{v}_{w}\right),$
where $t$ is the physical time, $H=\dot{a}/a$ is the Hubble parameter, and the momentum parameter $k_{w}=0.66\pm 0.04$ from the simulations Martins:2016ois . It is assumed in Eq. (85) that the curvature radii are identical to the correlation length $\overline{L}$. Taking the one-scale assumption $\overline{\rho}_{w}=\sigma_{w}/\overline{L}$ and adding the energy loss processes (damping effects) by hand to Eq. (85), one obtains the VOS model for the description of domain wall as Martins:2016book
$\frac{d\overline{L}}{dt}$ $=H\overline{L}+\frac{\overline{L}}{\ell_{d}}\overline{v}_{w}^{2}+c_{w}\overline{v}_{w},$ (86)
$\frac{d\overline{v}_{w}}{dt}$ $=\left(1-\overline{v}_{w}^{2}\right)\left(\frac{k_{w}}{\overline{L}}-\frac{\overline{v}_{w}}{\ell_{d}}\right),$
where the chopping parameter $c_{w}=0.81\pm 0.04$ from the simulations of domain wall evolution in the $Z_{2}$ model Martins:2016ois . The damping length is defined as
$\frac{1}{\ell_{d}}=3H+\frac{1}{\ell_{f}},$ (87)
which includes the damping effects on the network from Hubble drag and particle friction, which is related to the particle pressure $\Delta P$ by
$\frac{1}{\ell_{f}}=\frac{\Delta P}{\overline{v}_{w}\sigma_{w}}.$ (88)
Appendix D Beta function
In this section, we provide more information on the $\beta(x)$ function since it is critical to determine the PBH abundance. The $\beta(x)$ function is given by
$\beta(x)=\gamma_{p}{\rm erfc}(x),$ (89)
where the complementary error function is given by
[…]
The complementary error function represents the area under the two tails of a zero-mean Gaussian probability distribution with variance $\sigma^{2}=1/2$. The error function is related to the complementary error function via
${\rm erf}(x)=1-{\rm erfc}(x).$ (91)
Figure 10: Left: The $\beta(x)$ function as a function of $x$. Right: The value of $\beta$ as a function of the variance $\sigma$.
We plot $\beta(x)$ as a function of $x=\delta_{c}/(\sqrt{2}\sigma)$ in Fig. 10. We observe that to obtain an acceptable PBH abundance, $x$ should fall in the range of $3-5$ to exponentially suppress the PBH production. Therefore, the formation of PBHs via the critical density collapse mechanism arises from the tail of the probability density function.
Appendix E Strategy for PTA
In this section, we provide further details of the PTA data analysis. We follow the analysis procedure encoded in PTArcade, which wraps enterprise and Ceffyl to allow us to perform a standard search strategy for the PTA datasets (see Ref. Mitridate:2023oar for more details). The searches for PTA stochastic signals adopt the pulsars’ timing residuals, which are modeled as
$\overrightarrow{\delta t}=\vec{n}+\bm{F}\vec{a}+\bm{M}\vec{\epsilon},$ (92)
where the terms on the right-hand side of Eq. (92) are the white noise, red noise, and small errors in the fit to the timing-ephemeris parameter NANOGrav:2020tig , respectively.
Three types of white noise per backend/receiver system are included in the analysis, i.e., EFAC $E_{k}$, EQUAD $Q_{k}$[s], and ECORR $J_{k}$[s]. The pulsar-intrinsic red noise and GW background signals are included in the second term on the right-hand side of Eq. (92), which describes time-correlated stochastic processes. We adopt a power-law red noise with two parameters per pulsar for each dataset: the amplitude $A_{\rm red}$ and the spectral index $\gamma_{\rm red}$. For the GW signal, we adopt the domain wall annihilation scenario with four parameters in both datasets: the domain wall energy fraction $f_{w}(T_{\rm ann})$ at the annihilation temperature $T_{\rm ann}$, the domain wall annihilation temperature $T_{\rm ann}$ in units of GeV, the spectral index above the peak frequency $c_{1}$, and the parameter that represents the width of the spectral index $c_{2}$. The spectral $\gamma_{c}$ is fixed at 3 to respect the causality. In table 3, we provide the priors for the parameters of white and red noises, as well as the priors for parameters of the GW background signal from the annihilation of the domain walls.
In table 4 and Fig. 11, we report the posterior distributions for the parameters of the domain wall annihilations. The posteriors are obtained by GetDist. In comparison with the fit to the IPTA-DR2 data, the interpretation for NG15 data requires a larger domain wall energy density and a higher annihilation temperature. We also observe from Fig. 11 that the posterior distributions of $f_{w}$ and $T_{\rm ann}$ at $1\sigma$ and $2\sigma$ are centralized around the mean values while the posterior of the spectral parameters $c_{1}$ and $c_{2}$ have broad distributions, which may be due to large uncertainties in the data at high frequencies.
Table 3: The noise and signal parameters and their prior ranges.
Parameter Description Prior Comments
White noise
[…]
$Q_{k}$[s] EQUAD per backend/receiver system Log-uniform $[-8.5,\leavevmode\nobreak\ -5]$ Single-pulsar analysis only
$J_{k}$[s] ECORR per backend/receiver system Log-uniform $[-8.5,\leavevmode\nobreak\ -5]$ Single-pulsar analysis only
Red noise
$A_{\rm red}$ Red noise power-law amplitude Log-uniform $[-20,\leavevmode\nobreak\ -11]$ one parameter per pulsar
[…]
DW
$f_{w}(T_{\rm ann})$ Domain wall energy fraction at $T_{\rm ann}$ Log-Uniform $[-2,0]$ one parameter per PTA
[…]
Table 4: Bayes estimators for the fit to NG15 and IPTA-DR2 datasets.
NG15 IPTA-DR2
${\rm Log}_{10}f_{w}(T_{\rm ann})$ $-0.71\pm 0.17$ $-0.93\pm 0.10$
${\rm Log}_{10}[T_{\rm ann}/\rm GeV]$ $-0.70\pm 0.19$ $-1.28\pm 0.15$
$c_{1}$ 0.95$\pm$0.29 0.97$\pm$0.29
$c_{2}$ 1.66$\pm$0.55 1.48$\pm$0.58
Figure 11: Corner plot of the parameter posterior distributions with $1\sigma$ and $2\sigma$ contours.